OpenStudy (loser66):

A total x feet of fencing is to form 3 sides of a level rectangular yard. What is the maximum area of the yard in term of x? A) x^2/9 B) x^2/8 C) x^2/4 D)x^2 E)2x^2 Please, help

OpenStudy (freckles):

oops meant to write x

OpenStudy (freckles):

one sec

OpenStudy (freckles):

\[2L+W=x \\ A=LW=L(x-2L)\]

OpenStudy (freckles):

So we want to find A' here

OpenStudy (loser66):

yes, then??

OpenStudy (freckles):

x is a constant I believe \[A=Lx-2L^2 \\ A'=x-4L \\ A'=0 \implies x=4L \\ A=L(x-2L)=L(4L-2L)=L(2L)=2L^2 \\ \] I think you can find A in terms of x now since x=4L

OpenStudy (freckles):

still processing?

OpenStudy (loser66):

Yes. Thanks a lot. I got the right answer. It is not test, just practice exam with the answer sheet. But I want to know logic.

OpenStudy (freckles):

Why are you doing calculus 1 problems?

OpenStudy (loser66):

oh, is it calculus problem??

OpenStudy (freckles):

Yes it is an optimization problem

OpenStudy (loser66):

Question: why do you link to derivative?

OpenStudy (freckles):

because it say to maximize

OpenStudy (loser66):

I assume it is an algebra problem and get no outlet

OpenStudy (freckles):

usually those in the cal problems I have seen they actually gave a number instead of saying x

OpenStudy (loser66):

I am prepare for my GRE test for maths major students. ha!!!

OpenStudy (freckles):

http://tutorial.math.lamar.edu/Classes/CalcI/Optimization.aspx this example one here is pretty similar

OpenStudy (loser66):

My tough one is English, not math.

OpenStudy (freckles):

except as I said they say 500 feet instead of x feet

OpenStudy (loser66):

Any tips for GRE math practice?? I have only 1. :)

OpenStudy (freckles):

there was a question and I think it was on the gre it was my favorite question to give me a sec and I will remember it

OpenStudy (freckles):

solve for x+y \[x^2+y^2+2xy+2x+2y=-1 \]

OpenStudy (freckles):

there was a similar question on the gre or it might have been on another standard test can't remember which standard test

OpenStudy (loser66):

hihihi... I have time. I will study carefully before taking the test. But you know, I am old. hahaha..... I am slow!!

OpenStudy (freckles):

that question isn't too bad did you want me to show you how to do it or did you want to try it first

OpenStudy (loser66):

show me, please

OpenStudy (freckles):

\[x^2+2xy+y^2+2x+2y+1=0 \\ (x+y)^2+2(x+y)+1=0 \\ (x+y+1)^2=0 \\ x+y+1=0 \\ x+y=-1 \]

OpenStudy (freckles):

I turned it into a quadratic in terms of (x+y)

OpenStudy (loser66):

What if we let x+y =t??

OpenStudy (loser66):

oh, yea, we met there

OpenStudy (freckles):

good luck on your gre if i could figure out anything to tell you to help you I will I probably should have studied for it back in the day more than I did

OpenStudy (loser66):

Sure. Again, thanks for the help.

OpenStudy (freckles):

np

OpenStudy (loser66):

What is the units digit in standard decimal expansion of the number 7^(25) I don't get what it mean

OpenStudy (loser66):

since \(7^{25} = 1.34106862 *10^{21}\) hence the number of digits after decimal point is 8 but it is not one of the options which are 1,3,5,7 or 9

OpenStudy (loser66):

@freckles

OpenStudy (loser66):

To this problem, if I don't have calculator, how can I find out the answer?

OpenStudy (freckles):

so you are trying to find out the length of the number 7^(25) ?

OpenStudy (loser66):

problem 14 https://www.ets.org/s/gre/pdf/practice_book_math.pdf

OpenStudy (freckles):

on

OpenStudy (freckles):

oh*

OpenStudy (freckles):

the units digit is the ones digit

OpenStudy (freckles):

answer is 1

OpenStudy (loser66):

total??

OpenStudy (loser66):

nope, answer is 7

OpenStudy (freckles):

wait standard notation is not scientific notation is it ?

OpenStudy (loser66):

I am not sure.

OpenStudy (freckles):

there was need to write in scientific notation because standard notation is different from scientific notation

OpenStudy (freckles):

7^(25) mod 10 will give you number in 1's digit

OpenStudy (freckles):

or aka units digit

OpenStudy (freckles):

\[4^2=16 \text{; 16 is standard notation for } 4^2 \\ 4^2=1.6 \times 10 \text{ is sceinfict notation }\]

OpenStudy (freckles):

I missed the word no in there was need to write in scientific notation there was no need to write in scientific notation*

OpenStudy (loser66):

to this http://mathcentral.uregina.ca/QQ/database/QQ.09.06/tyrone1.html they said that the units digit is the last digit of the number in standard form in decimal form, the unit digit of 7^25 is the last one, which is 2, not 7 : (

OpenStudy (freckles):

\[7^{2} \mod 10=9 \\ 7^4 \mod 10 =9 \cdot 9 \mod 10=81 \mod 10=1 \\ 7^{4(6)+1} \mod 10=(7^4)^6(7) \mod 10=1^6(7) \mod 10 =7 \\ 7^{25} \mod 10=7 \]

OpenStudy (freckles):

the units digit is the first number it is the one's digit that is what that picture above says also

OpenStudy (freckles):

for example 61 do you call 1 the first digit because I do and 6 is the last digit 1 is in the one's place (or the units place) 6 is in the ten's palce

OpenStudy (freckles):

oh I think I see where you are getting confused

OpenStudy (freckles):

\[7^{25} = 1.34106862 *10^{21} \] you think this equality is exact when it is not

OpenStudy (freckles):

and even if it was exact you would say the units digit is 0 but it is not the case because this equation is not exact

OpenStudy (freckles):

like you still need to move that decimal to the right like 21 times

OpenStudy (freckles):

2 is not in the one's digit

OpenStudy (freckles):

http://www.wolframalpha.com/input/?i=7%5E25+

OpenStudy (freckles):

the number where it says result is the number in standard notation

OpenStudy (loser66):

oooooooooooooooooh!! got you

OpenStudy (freckles):

a mod 10 should always give you the units digit of a

OpenStudy (loser66):

modular method help me out 7^0 -----1 last digit 7^1 -----7 last digit 7^2 -----9 7^3 -----3 7^4-----1 hence the cycle is 4, 25 = 4*6 +1, hence \(7^25 = 7^24 *7 , but the last digit of 7^24 is 1 then *7 =7

OpenStudy (freckles):

yah

OpenStudy (loser66):

Thanks a ton. :) I do appreciate.

OpenStudy (freckles):

np

OpenStudy (freckles):

interesting practice exam I think they have made changes to it seen I last took the gre I feel like that was 8 years ago or something :p

OpenStudy (freckles):

I don't remember linear algebra or number theory on it

OpenStudy (freckles):

but maybe there was some